OpenStudy (anonymous):

What volume of chlorine gas at 42.0 deg C and 1.33 atm is required to react completely with 41.5 g of phosphorus according to the reaction: P4(s) + 6Cl2(g) -> 4PCl3(l)

OpenStudy (abmon98):

Number of moles=Mass(g)/Molar Mass(g/mol) you can look for phosphorus atomic weight in your periodic table P=30.974 30.974*4=123.896 g/mol Mass given: 41.5 g Number of moles of P4=41.5/123.896=0.334  Use Ideal gas Equation PV=nRT

OpenStudy (asib1214):

convert all given values into appropriate unit using appropriate conversion factors like convert 42C into Kelvins, ATM into kPa and grams into moles and plug those numbers into the following formula i'm going to provide you......